OpenStudy (anonymous):

Kevin sold 2 unicycles and 1 bicycle for a total of $110. Greg sold 4 unicycles and 3 bicycles for a total cost of $268. What is the cost of 1 bicycle?

OpenStudy (anonymous):

are there options?

OpenStudy (jziggy):

You have 2 equations and 2 unknowns, let unicycle = u and bicycle = b 2u + 1b = 110 4u + 3b = 268

OpenStudy (anonymous):

a. $27 b. $31 c. $48 d. $79

OpenStudy (jziggy):

using the first equation, b = 110 - 2u if you substitute that into the 2nd equation 4u + 3(110 - 2u) = 268 4u + 330 - 6u = 268 -2u = -62 u = 31 substitute that back into the first equation to get, b = 110 - 2(31) b = 48

OpenStudy (anonymous):

okay thank you! so like you had to combine the two equations then?

OpenStudy (jziggy):

Yes, I used the substitution method, which is to solve for one variable in terms of the other and then substitute it into the other equation giving you only one variable to work with.

OpenStudy (anonymous):

b=$48 type into wolfram like this: 2u+1b=110 and 4u+3b=268